10 millimeters are equal to _____ nanometers.

Question 9 options:

A)

103

B)

107

C)

10-3

D)

10-7
Answer #1

Some prefix used for subsidiary units

desi =      10 -1 m

centi =     10 -2 m        

milli   =      10 -3 m

micro =     10 -6 m

nano =      10 -9 m

pico =       10 -12 m

femto       10 -15 m

Answer

10 milli miter = ? Nano meter

1 millimiter = 10 -3 m

10 millimiter =10\times10 -3 m =10 -2 m

1 nano meter = 10 -9 m    Or 1 meter = 10 9 nano meter

1 meter      = 10 9 nano meter

10 -2 meter=10 -2 \times 10 9

                        =10 7 nano meter

10 millimiter =10 -2 m =10 7 nano meter

Answer is option B